OpenStudy (anonymous):

What is the sum of the whole number from 1-2000 please show work....

OpenStudy (anonymous):

1+2000 = 2001, 2+1999 =2001, 3+1998 = 2001, etc. there are 1000 pairs, so the sum is 1000*2001 = 2,001,000

OpenStudy (anonymous):

ohh..hey thanks!

OpenStudy (anonymous):

or Sum = (N/2)(N + 1) substituting, Sum = (2000/2)(2000+1) Sum = 2001000

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

oh..